# VecCity: A Taxonomy-guided Library for Map Entity Representation Learning [Experiment, Analysis & Benchmark]

### v1

@@ 4.4. Comparing of Modified Models
Review for MapRL. Pervious reviews (Atluri et al., 2018; Wang et al., 2022; Jin et al., 2023) discuss the applications of deep learning in urban data mining, providing a detailed overview of data, tasks, and deep learning models, particularly those that integrate sequence and graph encoders. However, their focus is on end-to-end models, while our work concentrates on pre-trained representation learning methods. Other studies (Zheng, 2015; Chen et al., 2024d) focus on data mining methods for trajectory data but do not cover pre-trained representation learning. Survey (Chen et al., 2024a) introduces a type-based taxonomy for MapRL where models are systematically reviewed based on the types of map entities they target. In contrast, our work does not classify methods by entity type but instead focuses on summarizing general techniques applicable to all map entities, i.e., a method-based taxonomy. This broader perspective helps uncover common design principles underlying MapRL methods.
Benchmark for MapRL. In recent years, the growing interest in MapRL has created a strong demand for open benchmarks to comprehensively analyze baseline models. Open-source benchmarking is valuable not only for advancing research but also for enabling practitioners to easily evaluate models and apply them to open datasets. While significant progress has been made in benchmarking for multivariate time series prediction (Wang et al., 2021; Shao et al., 2023; Qiu et al., 2024; Wang et al., 2024), computer vision (Kai Chen, 2019; Deng et al., 2009; Lin et al., 2014), and neural language processing (Pranav Rajpurkar, 2018; Conneau and Kiela, 2018; Reddy et al., 2019; Wang et al., 2019), benchmarking in the MapRL field remains largely unexplored. To the best of our knowledge, VecCity is the first benchmark specifically for MapRL, allowing researchers to conduct detailed comparative experiments and develop new models.
@@ 6. Conclusion
To overcome the challenges of fragmentation and the lack of standardized benchmarks in MapRL, we introduced a novel taxonomy organized by functional modules rather than entity types. Building on this taxonomy, we developed VecCity, a modular library offering interfaces for encoding, pre-training, fine-tuning, and evaluation. Using VecCity, we established the first standardized benchmarks by reproducing 21 mainstream models and integrating datasets from nine cities. Moreover, we also conduct an in-depth evaluation of mainstream MapRL models. Our comprehensive experiments highlight the impact of pre-training tasks and encoder architectures, demonstrating the advantages of combining multiple components. VecCity provides a unified framework that promotes reusability, streamlines experimentation, and advances research in MapRL. Although VecCity currently supports most mainstream MapRL models, we plan to expand it further to cover a wider range of pre-trained spatiotemporal data representation learning algorithms, such as trajectory representation learning.

### v2

@@ 4.3.2. Efficiency
Niche Tasks for MapRL. Raster and vector data models are the two primary formats for electronic maps. The raster model structures spatial information through grids, which are commonly used in remote sensing and environmental analysis. In contrast, the vector model represents map entities as points, polylines, and polygons, leveraging graph structures to represent relations between entities. These structural differences lead to different modeling approaches: CNNs are usually used for raster-based learning, while GNNs better capture spatial relations in vector data. As a result, raster and vector models have distinct benchmarks. Due to its high precision and accuracy, the vector data model serves as the fundamental representation for electronic maps.
In parallel, various downstream tasks in MapRL have been widely explored, including Next POI Prediction (NPP), Travel Time Estimation (TTE), and Mobility Inference (MI). These tasks play a key role in location-based services and spatiotemporal data mining. For NPP, models like DeepMove(Feng et al., 2018) employ hierarchical RNNs with attention mechanisms to capture sequential patterns in user trajectories. TTE methods such as DeepTTE (Wang et al., 2018) integrate LSTM networks with dynamic graph convolutional layers, fusing traffic state and weather data for accurate time estimation. For MI, models like ODMP (Wang et al., 2019) adopt a full-pipeline approach, integrating sequence and graph encoders to effectively capture demand-supply dynamics in ride-hailing scenarios.
Reviews for MapRL. Pervious reviews (Atluri et al., 2018; Wang et al., 2022; Jin et al., 2023) discuss the applications of deep learning in urban data mining, providing a detailed overview of data, tasks, and deep learning models, particularly those that integrate sequence and graph encoders. However, their focus is on end-to-end models, while our work concentrates on pre-trained representation learning methods. Other studies (Zheng, 2015; Chen et al., 2024d) focus on data mining methods for trajectory data but do not cover pre-trained representation learning. Survey (Chen et al., 2024a) introduces a type-based taxonomy for MapRL where models are systematically reviewed based on the types of map entities they target. In contrast, our work does not classify methods by entity type but instead focuses on summarizing general techniques applicable to all map entities, i.e., a method-based taxonomy. This broader perspective helps uncover common design principles underlying MapRL methods.
Benchmark for MapRL. The growing interest in MapRL has increased the demand for open benchmarks to analyze baseline models. Open-source benchmarking is valuable not only for advancing research but also for enabling users to evaluate models and apply them to open datasets. While significant progress has been made in benchmarking for multivariate time series prediction (Wang et al., 2021; Qiu et al., 2024; Wang et al., 2024), computer vision (Deng et al., 2009), and neural language processing (Pranav Rajpurkar, 2018), benchmarking in the MapRL field remains largely unexplored. To our knowledge, VecCity is the first MapRL benchmark for comparative experiments and model development.
@@ 7. Conclusion
To overcome the challenges of fragmentation and the lack of standardized benchmarks in MapRL, we introduced a novel taxonomy organized by functional modules rather than entity types. Building on this taxonomy, we developed VecCity, a modular library offering interfaces for encoding, pre-training, fine-tuning, and evaluation. Using VecCity, we established the first standardized benchmarks by reproducing 21 mainstream models and integrating datasets from nine cities. Moreover, we also conduct an in-depth evaluation of mainstream MapRL models. Our comprehensive experiments highlight the impact of pre-training tasks and encoder architectures, demonstrating the advantages of combining multiple components. VecCity provides a unified framework that promotes reusability, streamlines experimentation, and advances research in MapRL. Although VecCity currently supports most mainstream MapRL models, we plan to expand it further to more pre-trained spatiotemporal data representation learning algorithms.